# Cosmic topology. Part Ic. Limits on lens spaces from circle searches

### v1

@@ 1 Introduction
This search for matched-circle pairs was performed in full generality on the Wilkinson Microwave Anisotropy Probe (WMAP) full-sky temperature map and no statistically significant matches were found [14, 22, 23]. The search was repeated on Planck 2013 maps with identical results [24], but not reported. A limited search was conducted by the Planck team, again with negative outcome [16, 17]. In all cases, the smallest circle that could be ruled out depended on the required false negative and positive rates. However, both are steep functions of the radius of the circle for small circle size.444 A more subtle question is how well to trust limits on those circles that are small enough to lie entirely or mostly within the usual foreground masks that are applied to full-sky CMB temperature maps. We reserve such questions for a future paper revisiting the statistical details of the circle searches.
@@ 5 Cosmological constraints on detectability of lens spaces
We note that the parameter space allowed by the lack of matched circles, although much better constrained, is still large. These spaces are inhomogeneous for $q>1$ and could therefore be observed or constrained by their effect on cosmological observables such as the CMB anisotropies; we will study their exact effect in an upcoming paper.
As noted in the introduction, observational constraints on lens spaces had been considered in Ref. [27] prior to searches for circles in the sky. When applied to the LSS, that work provides an alternative and $q$-dependent upper bound on allowed lens spaces. They required
$\frac{2\pi q}{p}<f_{\mathrm{O}}d_{\mathrm{LSS}}\,,$ (5.4)
which translates into the upper bound
@@
The tightest constraint on lens spaces is determined from a combination of these two bounds. No observer will see circles when

### v2

@@ 1 Introduction
This search for matched-circle pairs was performed in full generality on the Wilkinson Microwave Anisotropy Probe (WMAP) full-sky temperature map and no statistically significant matches were found [14, 22, 23]. The search was repeated on Planck 2013 maps with identical results [24], but not reported. A limited search was conducted by the Planck team, again with negative outcome [16, 17]. In all cases, the smallest circle that could be ruled out depended on the required false negative and positive rates. However, both are steep functions of the radius of the circle for small circle size.444 A more subtle question is how well to trust limits on those circles that are small enough to lie entirely or mostly within the usual foreground masks that are applied to full-sky CMB temperature maps. We reserve such questions for a future paper revisiting the statistical details of the circle searches. If the topology scale of the Universe were sufficiently small to produce matched circles on the CMB sky, they would appear in both temperature and polarization. Therefore, future full-sky observations of the CMB polarization are expected to provide a complementary verification of the lack of circles; this probe is valuable as it is sensitive to different systematics and foregrounds, but it has limited discovery power as CMB temperature already constrains most of the parameter space accessible by circle searches.
@@ 5 Cosmological constraints on detectability of lens spaces
We note that the $(p,q)$ parameter space allowed by the lack of matched circles is still large given that our current knowledge of $\Omega_{K}$ is poor. In particular, the closer $\Omega_{K}$ is to zero, the weaker are the constraints on $p$ and $q$. Of course, there is more cosmological information available in the CMB than just whether or not there are matched circles; we will study the broader effects of topology on CMB anisotropies in $S^{3}$ with non-trivial topology in future papers.
As noted in the introduction, observational constraints on lens spaces had been considered in Ref. [27], where the authors demonstrated that, for a given $(p,q)$, the absolute (“global”) maximum distance to the nearest clone is $2\pi q/p$ in units of $R_{c}$. (This is realized at $s=0$ for the $j=1$ clone, as seen from Eq. 3.1.) When applied to the LSS, and given the non-observation of matched circles, this provides an alternate $q$-dependent upper bound on allowed lens spaces, requiring
$\frac{2\pi q}{p}<f_{\mathrm{O}}d_{\mathrm{LSS}}\,.$ (5.4)
This translates into the upper bound
@@
The tightest constraint on lens spaces is determined from a combination of the two bounds (5.3) and (5.5): no observer will see circles when